what is the y-intercept and x-intercept for the equation -6x+y=26
Answer by Alan3354(69443)

To find the x-int, set y = 0.
-6x+y=26
-6x = 26
x = -13/3
------------
To find the y-int, set x = 0.
-6x+y=26
y = 26
--------